# Biomedical Image Segmentation: A Systematic Literature Review of Deep Learning Based Object Detection Methods

### v1

@@ 4.2 One-stage Detection Models
1) You Only Look Once (YOLO): In the two-stage object detection algorithms e.g., Fast R-CNN, the model first identifies the regions of interest with the help of RPN and in the second stage each region of interest is given class labels for detection. In case of YOLO, both these functions are performed simultaneously with a fully connected convolutional layer. After YOLO was introduced in 2015, so far, few versions have been developed each having different pros and cons. The current variants of YOLO include YOLOV2 to YOLOV8. The working mechanism of YOLO consists of few stages as shown in  [47].
2) Single Shot Detector (SSD): SSD is a one-stage object detection model consisting of two parts namely the backbone and the SSD head. The backbone is a type of pre-trained convolutional neural network used for feature extraction. The SSD head is one convolutional layer of few convolutional layers that generates the bounding box and class of the object. The complete working procedure of SSD is shown in  [48].
@@
4) RetinaNet: It is another one-stage detection model consisting of one backbone and two subnetworks. The backbone is a convolutional neural network like ResNet used for feature map generation of the whole image. The first subnetwork performs the classification based on the outputs of backbone model whereas the second subnetwork is responsible for identification of bounding boxes and the class of each bounding box  [50].
@@ 4.3 Point Based Detection Models
There are some object detection algorithms which can applied in procedures sightly different both from one-stage detection techniques as well as two-stage detection techniques. All these algorithms have been placed in a separate category. Their working mechanisms and power are higher than both the already discussed two-stage detectors as well as one-stage detectors. Therefore, they have been categorized as point based detectors because they use specific points for detection of objects in inages. Some of these detectors are discussed in this section
2) CornerNet: In cornerNet, the bounding boxes of objects inside image are detected using pair key points namely top-left corner and bottom right corner. This process of finding keypoints is performed by using convolutional neural network. By using this procedure, the need for designing anchor boxes which are applied in previously discussed one-stage detecting approaches is eliminated  [51].
@@
Foveabox: FoveaBox is flexible, accurate, and anchor free architecture for object detection. All the previous anchor-based object detection models utilize anchors to detect objects inside an image, their performance and ability of generalization are limited to the design of anchors. In contrast, the Foveabox directly identifies boundary boxes and the existence of objects in the image without refereeing to the pre-defined anchors  [54].
5 SELECTED ARTICLES AND THEIR ANALYSIS
@@ 5.1 Articles related to Two-stage Detection
Figure 5: A chart showing the modalities (light green color) and diseases (yellow color). The numeric values indicates the number of papers with each modality.
The authors in  [79] applied mask R-CNN for the segmentation of ultrasound images. They introduced some modification in the RPN for improving the segmentation accuracy of the mask R-CNN model. Similarly, a combination of mask R-CNN and 3D U-Net was applied by the researchers in  [80] for the segmentation of CT images of pancreas. The authors in  [81] used mask R-CNN for segmentation of lungs nodule. The proposed model consists of backbone network, RPN, ROI Align and fully convolutional neural network. The authors in  [82] proposed a combination of mask R-CNN and holistically nested edge detection (HED) model for the segmentation of MRI and CT images of liver.
Similarly, mask R-CNN was used by Linqin et al for detection and segmentation to be used applied for 3D visualization of pulmonary nodule  [83]. Zhiyong et al.  [84] suggested a hybrid model of mask R-CNN and inception-V3 for the segmentation of ultrasound images to detect prostate cancer. In the proposed model sobel convolution has been applied for sharpening the input images. Yuanhao et al.  [85] performed the classification and segmentation of ultrasound images for detection of breast tumour. They applied superpixel elastic deformation approach for data augmentation and mask R-CNN for segmentation of the ultrasound images for possible identification of breast tumour.
Similarly, Jui-ying et al.  [86] proposed mask R-CNN for the segmentation of sonogram images to detect breast cancer. The architecture they developed consists of CNN for feature map generation, RPN for extracting regions of interest, ROI align for extracting features of each ROI. In the proposed model, the RPN has been combined with fast R-CNN. For the assessment of fetal growth in ultrasound images, Moccia et al.  [87] suggested an advanced version of mask R-CNN known as mask R2-CNN. Meng et al.  [88] suggested mask R-CNN for the segmentation and detection of adenomatous colorectal polyps in the colonoscopy images. Similarly, Jin et al.  [89] used mask R-CNN for the detection of early gastric cancer in the endoscopy images. Qadir et al.  [90] proposed mask R-CNN for the detection and segmentation of polyp in the colonoscopy images. Jung et al  [91] applied mask R-CNN for the automatic segmentation of nuclei in histopathology images. Santos et al.  [92] applied the mask R-CNN for the segmentation of fundus images to detect lesions in diabetic retinopathy images.
Table 2: TWO STAGE DETECTORS FOR BIOMEDICAL IMAGING SEGMENTATION
@@
Table 2 (Continued) ID-Ref Year Modality Model Disease $||$ Organ Evaluation Metrics Results A042  [96] 2021 PET CT Mask R-CNN Tumour Prostrate ACD, VD, DSC 0.83+-0.91, -0.01+-+-0.79, 0.84+-0.09 A043  [97] 2024 Microscopic Images Mask R-CNN Cell Detection Dice, Accuracy, Precision, Sensitivity, F1 Score, Specificity, Volume Difference 0.862, 0.945, 0.901, 0.827, 0.862, 0.977, 0.082 A044  [98] 2023 Microscopic Images Mask R-CNN Anthrax Tissue Disease AP, AP75, IoU, DICE 0.560, 0.613, 0.751, 0.840 A045  [99] 2021 Microscopic Images Mask R-CNN Cell Segmentation SEG, DET, TRA, D-T 85.26%, 97.14%, 97.05%, 0.09 A046  [100] 2023 Ultrasound Cascade R-CNN, Faster R-CNN Arteries Disease Accuracy, AO, NRA, RASIRA 86.5%, 90.4%, 84.7, 88.8% A047  [101] 2022 Ultrasound Cascade R-CNN Thyroid Nodules mAP, TPR, FPR, TNR, FNR, Accuracy 87.1%, 0.9841, 0.0082, 0.9894, 0.0158, 0.9867 A048  [102] 2022 Ultrasound Cascade R-CNN Breast Cancer Accuracy, Jaccard, Precision, Recall, Specificity 97.51%, 72.47%, 82.70%, 99.13%, 79.35% A049  [103] 2019 Endoscopy Cascade R-CNN FPN Multi-class Artifacts Score Detection, Score Segmentation 0.3429, 0.3500 A050  [104] 2019 Endoscopy Cascade R-CNN Artifacts mAP, IoU, Score 0.3235, 0.4172, 0.3610 A051  [105] 2023 Ultrasound FPN Fetal head segmentation Heart DICE, IoU, Hausdorff, Conformity 89.08%, 80.70%, 19.76, 74.64% A052  [106] 2023 Endoscopy FPN Polyp Different Varying different varying A053  [107] 2021 Endoscopy FPN + CNN Polyp Average Precision, Segmentation Score 0.8986, 0.7771 A054  [108] 2015 Histopathology FCNN Breast Regions Average Precision, Average Recall, Average F1 Score, Variance Precision, Variance Recall, Variance F1 Score 0.91, 0.82, 0.85, 0.015, 0.02, 0.01 A055  [109] 2019 Fundus FCNN Diabetic Retinopathy DSC, Sensitivity, Specificity, Accuracy 96.7%, 98.1%, 99.9%, 99.9% A056  [110] 2018 Fundus FCNN Vessels Detection Accuracy, Sensitivity, Specificity 0.9210, 0.7215, 0.9576 A057  [111] 2017 Fundus FCNN Retinal Segmentation Sensitivity, Specificity, PPV, F1 Score 93.12%, 99.56%, 89.90%, 90.93% A058  [112] 2018 Biopsy FCN Dermis Epidermis RCPC, Jaccard 0.8801, 0.8845, 0.8900, 0.9962 A059  [113] 2018 PET CT FCN Tumour DSC, Classification Error, Volume Error 0.85, 0.33, 0.15 A060  [114] 2024 PET FCN Lymphoma Detection Sensibility, Dice 0.9953, 0.8665 A061  [115] 2020 Dermoscopy Colonoscopy Endoscopy U-Net Multiple mIoU, DSC, Recall, Precision 0.6255, 0.7649, 0.7156, 0.8007 A062  [116] 2023 Sonograms D-CNN U-Net Breast Cancer Accuracy, Sensitivity, Specificity, AUC, UIoU, MDice 97%, 97.7%, 96.4%, 0.96, 0.89, 0.92 A063  [117] 2019 Angiographic CT MRI CE-Net Lungs, Vessels, Eyes Sensitivity, Accuracy, AUC 0.8309, 0.9545, 0.9779 A064  [118] 2021 Colonoscopy ColonSegNet Polyp AP, IoU, DSC, Jaccard, Precision, Recall 0.8000, 0.8100, 0.7239, 0.8435, 0.8496, 0.9493
Huang et al.  [93] applied mask R-CNN for the segmentation of dermatology images to detect lesions in the skin. Similarly, Bagheri et al.  [94] suggested the application of mask R-CNN, Retina-Deeplab, and graph-based model for the segmentation of dermoscopic images to detect lesions in the skin. Zhang et al.  [95] applied mask R-CNN for the segmentation of PET images to detection tumour in human lungs. Similarly, Wang et al.  [96] applied dual mask R-CNN for the segmentation of PET and CT images to detect tumour and prostrate. Wang et al.  [97] applied mask R-CNN for the analysis of microscopic images. Similarly, Hoorali et al.  [98] proposed an enhanced version of mask R-CNN called URCNN for the segmentation of microscopic images to detect anthrax and some other tissue diseases. Similarly, Chen et al.  [99] suggested the use of faster R-CNN for the segmentation of microscopic images to detect cancer cells. The authors have applied faster R-CNN and the circle scanning algorithm for improving the detection of cancer cells.
@@
Wang et al.  [100] tested three two-stages model namely cascade R-CNN, faster R-CNN, and double head R-CNN on ultrasound images of arteries for their segmentation to detect various abnormalities. Zheng et al.  [101] introduced and improved version of cascade R-CNN for the segmentation of ultrasound images to detect thyroid nodules. Chen et al.  [102] proposed cascade R-CNN for the segmentation of breast ultrasound images to detect lesions. Zhang and Xie  [103] suggested a hybrid model of FPN and cascade R-CNN for the detection and segmentation of different multi-class artifacts in the endoscopy images. Yang and Cheng  [104] suggested the implementation of cascade R-CNN for the detection and segmentation of artefacts in the endoscopic images. In another work, Chen et al.  [105] proposed a modified version of feature pyramid network (FPN) called Fourier feature pyramid network (FFPN) for the segmentation of ultrasound images. Su et al.  [106] proposed an improved version of FPN called FedNet for the segmentation of polyp in the endoscopic images. Wuyang et al.  [107] suggested FPN with other CNN networks for the detection and segmentation of polyp in the endoscopy images. Su et al.  [108], suggested an improved version of CNN named fCNN standing for fast scanning deep convolutional neural network for the segmentation and identification of regions in histopathology images of breast. Xue et al.  [109] proposed a multi-task segmentation approach for the segmentation of retina for the detection of diabetic retinopathy. A fully connected convolutional neural network was proposed by Li et al.  [110] for the segmentation of retinal fundus images to detect vessels in the images. Similarly, the fully connected convolutional neural network was applied by Feng et al.  [111] for the segmentation of retinal images. Pal et al.  [112] applied deep convolutional neural network (FCN) for the segmentation of biopsy images to detect dermis, epidermis, and non-tissue regions. Xhao et al.  [113] proposed a fully connected convolutional neural network for the segmentation of PET and CT images to detect tumour.
@@
Pang et al.  [114] suggested CNN for the segmentation of PET images based on multi-temporal approach to detect lymphoma. For improving the performance of standard U-Net architecture, the authors in  [115] developed an improved version of the standard U-Net called double U-Net in which they stacked two U-Nets on the top of each other. The model was evaluated with four publicly available dataset of different modalities e.g., dermoscopy, colonoscopy, and microscopy. Yang et al.  [116] suggested the use of attention U-Net and deep convolutional neural network for the segmentation and classification of sonography ultrasound images for the detection of breast cancer. A model called CE-Net standing for context encoder network was applied for the segmentation of 2D medical imaging in the study presented in  [117]. JHA et al.  [118] proposed ColonSegnet for the detection, localization, and segmentation of polyp in the colonoscopy images.
@@ 5.2 Articles related to One-stage Detection
One-stage object detection models for the segmentation of biomedical imaging include, YOLO, SSD, YOLOv1/v2/v3/v4, RefineNet, and RetinaNet. In this section, these models applied for the segmentation of biomedical images segmentation are discussed in detail along with other considerations. For example, Kavitha and Palaniappan  [119] applied YOLO for the segmentation of brain MRI images to detect tumour. Similarly, Chang et al.  [120] suggested two deep learning models being YOLO and fully connected convolutional neural network for the segmentation and classification of heart MRI images. Suryawanshi and Patil  [121] applied YOLOv5 for the segmentation of MRI images to detect brain tumour. Similarly, Moreno et al.  [122] applied a combination of two convolutional neural networks for the segmentation of left ventricle in heart MRI images. In the same fashion, Zhong and Deng  [123] introduced an improved version of YOLOv3 for the segmentation of MRI images for the detection of herniated intervertebral discs. Unver and Ayan  [124] used YOLO and Grabcut model algorithm for the segmentation of dermoscopic images to detect lesions in the skin. Similarly, Lemay  [125] applied YOLOv3 for the segmentation of 2D and 3D CT images for localizing the kidneys in the images and diseases including Cystic kidneys, hypertrophied kidneys, and tumoral kidneys.
Pang et al.  [126] suggested an enhanced model of YOLOv3 called YOLOv3-arch for the segmentation of CT images to identify cholelithiasis and classify gallstones. Similarly, Wang et al.  [127] applied the YOLOv2 and Fourier Photography Microscopy (FPM) in combination for the detection of WBC. Afshari et al.  [128] used YOLO to localize and detect different normal active organs in the 3D PET imaging modality. Similarly, Nguyen et al.  [129] used Faster R-CNN and YOLO for the segmentation of chest X-Ray images to detect different abnormalities. In a similar attempt, Liu et al.  [130] proposed the YOLOv8 for the X-Ray angiography processing to segment the coronary artery. Wu et al.  [131] used YOLOv5 for multi-type lesions detection in angiography images. Xu et al.  [132] used scaled-YOLOv4 for the detection of pulmonary embolism in the CT angiograms. Similarly, Mortada et al.  [133] applied YOLOv7 for the segmentation of echocardiographic images to identify and segment the anatomical structures of left heart. Pkhrel et al.  [134] suggested the use of YOLOv8 along with the augmentation for the segmentation of coronary artery in the angiograms to detect diseases. In another study, Osama et al.  [135] used YOLOv8 model for the segmentation of angiography images to identify and detect coronary artery stenosis. Zhuang et al.  [136] used YOLOv3 for the segmentation of left ventricle in the echocardiography. Similarly, Zhang et al.  [137] suggested the use of and enhanced model YOLOV3 one-stage segmentation model for segmentation of ultrasound images to detect the location of thyroid nodules.
@@
Lu et al.  [138] used YOLOX for the segmentation of ultrasound images to examine the fetus for congenital heart disease. And enhanced model of YOLOV5 called YOLO-CSE was proposed by Hussein et al.  [139] for the segmentation of muscle ultrasound to detect myositis. Xie et al.  [140] presented YOLOV4 in their work for the segmentation of ultrasound images to detect different organs. In the same manner, Ghose et al.  [141] suggested the use of an enhanced version of YOLOV5 to segment the colonoscopic images for the detection of polyp. Doniorjon et al.  [142] used YOLOV4-tiny for the segmentation of endoscopic images to detect polyp. Kirthika and Sargunam  [143] applied YOLOV4 for the segmentation of endoscopic images to detect multi-class artefact. Durak et al.  [144] implemented different types of deep learning models for the detection of polyp in the gastro endoscopic images. Al Zorgani et al.  [145] proposed YOLOV2 for the detection of breast cancer in histopathological breast images. Rong et al.  [35] applied YOLO for the segmentation of nucleus and characterization of tumour microenvironment. Ayaz et al.  [146] proposed different deep learning models including mask R-CNN, U-Net, CNN, and YOLO for the segmentation of nucleus in the histopathological images. Narotamo et al.  [147] applied YOLO for detection of nucleus and U-Net for the segmentation of nucleus in fluorescence microscopy. Nair et al.  [148] applied YOLOV4 for the nucleus detection in the histopathology images of breast.
@@
Similarly, Bsl et al.  [149] suggested YOLO for the identification of region proposal network to diagnose cancer in the breast histopathology images. Similarly, Pal et al.  [150] applied YOLOV3 for detection of lesions in the retinal fundus images. Lalitha and Padyana  [151] in fundus images using YOLO-RF which is a hybrid model of YOLO and random forest. Santos et al.  [152] used YOLOV5 for lesion detection in fundus images. Saha et al.  [153] implemented CNN with YOLO for the detection of glaucoma in the color fundus images. Ramachandran et al.  [154] proposed the application of YOLOv3 for the localization of optic disc in retinopathy fundus images. Bagheri et al.  [155] applied YOLO for lesion detection in skin using dermoscopic images. Singh et al.  [156] suggested fuzzy logic in combination with object detection models for the detection of cancer in the dermoscopic images. Yin  [157] applied YOLOV3 network for the segmentation of PET/CT to detect tumour in Lymphoma patients. Shwetha et al.  [158] performed a two-stages approach for the detection of pus cell and epithelial cell. Similarly, Kong and Shen  [159] proposed an enhanced version of YOLO for the segmentation of microscopic images to detect microorganisms.
@@
Puerta et al.  [160] applied a combined model of YOLO and U-Net for the segmentation of microscopic images to characterize the neuron. Cao et al.  [161] suggested SSD and other segmentation models for the detection of benign and malignant breast cancer in the ultrasound images. Similarly, the SSD was applied by Kulhare et al.  [162] for the detection of different lungs abnormalities in ultrasound images. Ansari  [163] proposed a modified form of single shot detector named MP-FSSD (Multi-scale Pyramidal Fusion Single Shot Multibox Detector Network) for the segmentation of endoscopic images frames or the endoscopy to detect polyp. Chen et al.  [164] proposed SSD for the detection of Polyp in the gastrointestinal endoscopic images using the segmentation. In a similar attempt, Xu et al.  [165] suggested DAD-SSD (Dynamic Attention Deconvolutional Single Shot Detector) for the classification of different types of polyps in the colonoscopic images. De Lucena et al.  [166] applied RetinaNet for the detection of skin different diseases namely melanoma, dysplastic nevus, and healthy skin. Rehman et al.  [167] suggested deep RetinaNet for the segmentation of dermoscopic images to identify the melanoma lesion. A hybrid model of RetinaNet and ResNet was applied for the detection of skin cancer in dermoscopic images by Ahmed et al.  [168]. Bagheri et al.  [169] proposed three different object detection models for the identification of skin lesions in the dermoscopic images.

### v2

@@ 4.2 One-stage Detection Models
1) You Only Look Once (YOLO): In the two-stage object detection algorithms e.g., Fast R-CNN, the model first identifies the regions of interest with the help of RPN and in the second stage each region of interest is given class labels for detection. In case of YOLO, both these functions are performed simultaneously with a fully connected convolutional layer. After YOLO was introduced in 2015, so far, few versions have been developed each having different pros and cons. The current variants of YOLO include YOLOV2 to YOLOV8. The working mechanism of YOLO consists of few stages [51, 52, 53, 54, 55, 56, 57].
2) Single Shot Detector (SSD): SSD is a one-stage object detection model consisting of two parts namely the backbone and the SSD head. The backbone is a type of pre-trained convolutional neural network used for feature extraction. The SSD head is one convolutional layer of few convolutional layers that generates the bounding box and class of the object. The complete working procedure of SSD is shown in  [58, 59, 60, 61, 62, 63].
@@
4) RetinaNet: It is another one-stage detection model consisting of one backbone and two subnetworks. The backbone is a convolutional neural network like ResNet used for feature map generation of the whole image. The first subnetwork performs the classification based on the outputs of backbone model whereas the second subnetwork is responsible for identification of bounding boxes and the class of each bounding box [65, 64, 66, 67, 68].
@@ 4.3 Point Based Detection Models
There are some object detection algorithms which can applied in procedures sightly different both from one-stage detection techniques as well as two-stage detection techniques. All these algorithms have been placed in a separate category. Their working mechanisms and power are higher than both the already discussed two-stage detectors as well as one-stage detectors. Therefore, they have been categorized as point based detectors because they use specific points for detection of objects in images. Some of these detectors are discussed in this section
1) CornerNet: In cornerNet, the bounding boxes of objects inside image are detected using pair key points namely top-left corner and bottom right corner. This process of finding keypoints is performed by using convolutional neural network. By using this procedure, the need for designing anchor boxes which are applied in previously discussed one-stage detecting approaches is eliminated  [69].
@@
4) Foveabox: FoveaBox is flexible, accurate, and anchor free architecture for object detection. All the previous anchor-based object detection models utilize anchors to detect objects inside an image, their performance and ability of generalization are limited to the design of anchors. In contrast, the Foveabox directly identifies boundary boxes and the existence of objects in the image without refereeing to the pre-defined anchors  [72].
5 SELECTED ARTICLES AND ANALYSIS
@@ 5.1 Articles related to Two-stage Detection
Figure 5: A chart showing the modalities (light green) and diseases (yellow). The numeric values indicates the number of papers with each modality.
Liu and Li [97] applied mask R-CNN for the segmentation of ultrasound images. They introduced some modification in the RPN for improving the segmentation accuracy of the mask R-CNN model. Similarly, a combination of mask R-CNN and 3D U-Net is used for the segmentation of CT images of pancreas [98]. Furthermore, mask R-CNN has also been used for segmentation of lungs nodule [99], and 3D visualization of pulmonary nodule  [100]. The proposed model [99] consists of backbone network, RPN, ROI Align and fully convolutional neural network. The authors in  [101] proposed a combination of mask R-CNN and holistically nested edge detection (HED) model for the segmentation of MRI and CT images of liver.
Recently, Zhiyong et al.  [102] proposed a hybrid model of mask R-CNN and inception-V3 for the segmentation of ultrasound images to detect prostate cancer. In the proposed model sobel convolution has been applied for sharpening the input images. Similarly, Yuanhao et al. [103] performed the classification and segmentation of ultrasound images for detection of breast tumour. They applied superpixel elastic deformation approach for data augmentation and mask R-CNN for segmentation of the ultrasound images for possible identification of breast tumour. Another study used mask R-CNN architecture for the segmentation of sonogram images to detect breast cancer [104]. This architecture [104] includes a CNN for feature map generation, an RPN for extracting regions of interest, and ROI Align for extracting features of each ROI. In the proposed model, the RPN has been combined with fast R-CNN.
For assessing fetal growth in ultrasound images, an advanced version of mask R-CNN, known as Mask R2-CNN [105] is used. Furthermore, mask R-CNN has been employed for the segmentation and detection of adenomatous colorectal polyps in colonoscopy images [106], the detection of early gastric cancer in endoscopy images [107], the automatic segmentation of nuclei in histopathology images [108], and the segmentation of fundus images to detect lesions in diabetic retinopathy [109].
Table 2: Two-Stage Detectors for Biomedical Image Segmentation.
@@
A042  [114] 2021 PET CT Mask R-CNN Tumour Prostrate ACD, VD, DSC 0.83+-0.91, -0.01+-+-0.79, 0.84+-0.09
A043  [115] 2024 Microscopic Images Mask R-CNN Cell Detection Dice, Accuracy, Precision, Sensitivity, F1 Score, Specificity, Volume Difference 0.862, 0.945, 0.901, 0.827, 0.862, 0.977, 0.082
A044  [116] 2023 Microscopic Images Mask R-CNN Anthrax Tissue Disease AP, AP75, IoU, DICE 0.560, 0.613, 0.751, 0.840
A045  [117] 2021 Microscopic Images Mask R-CNN Cell Segmentation SEG, DET, TRA, D-T 85.26%, 97.14%, 97.05%, 0.09
A046  [118] 2023 Ultrasound Cascade R-CNN, Faster R-CNN Arteries Disease Accuracy, AO, NRA, RASIRA 86.5%, 90.4%, 84.7, 88.8%
A047  [119] 2022 Ultrasound Cascade R-CNN Thyroid Nodules mAP, TPR, FPR, TNR, FNR, Accuracy 87.1%, 0.9841, 0.0082, 0.9894, 0.0158, 0.9867
A048  [120] 2022 Ultrasound Cascade R-CNN Breast Cancer Accuracy, Jaccard, Precision, Recall, Specificity 97.51%, 72.47%, 82.70%, 99.13%, 79.35%
A049  [121] 2019 Endoscopy Cascade R-CNN FPN Multi-class Artifacts Score Detection, Score Segmentation 0.3429, 0.3500
A050  [122] 2019 Endoscopy Cascade R-CNN Artifacts mAP, IoU, Score 0.3235, 0.4172, 0.3610
A051  [123] 2023 Ultrasound FPN Fetal head segmentation Heart DICE, IoU, Hausdorff, Conformity 89.08%, 80.70%, 19.76, 74.64%
A052  [124] 2023 Endoscopy FPN Polyp Different Varying different varying
A053  [125] 2021 Endoscopy FPN + CNN Polyp Average Precision, Segmentation Score 0.8986, 0.7771
A054  [126] 2015 Histopathology FCNN Breast Regions Average Precision, Average Recall, Average F1 Score, Variance Precision, Variance Recall, Variance F1 Score 0.91, 0.82, 0.85, 0.015, 0.02, 0.01
A055  [127] 2019 Fundus FCNN Diabetic Retinopathy DSC, Sensitivity, Specificity, Accuracy 96.7%, 98.1%, 99.9%, 99.9%
A056  [128] 2018 Fundus FCNN Vessels Detection Accuracy, Sensitivity, Specificity 0.9210, 0.7215, 0.9576
A057  [129] 2017 Fundus FCNN Retinal Segmentation Sensitivity, Specificity, PPV, F1 Score 93.12%, 99.56%, 89.90%, 90.93%
A058  [130] 2018 Biopsy FCN Dermis Epidermis RCPC, Jaccard 0.8801, 0.8845, 0.8900, 0.9962
A059  [131] 2018 PET CT FCN Tumour DSC, Classification Error, Volume Error 0.85, 0.33, 0.15
A060  [132] 2024 PET FCN Lymphoma Detection Sensibility, Dice 0.9953, 0.8665
A061  [133] 2020 Dermoscopy Colonoscopy Endoscopy U-Net Multiple mIoU, DSC, Recall, Precision 0.6255, 0.7649, 0.7156, 0.8007
A062  [134] 2023 Sonograms D-CNN U-Net Breast Cancer Accuracy, Sensitivity, Specificity, AUC, UIoU, MDice 97%, 97.7%, 96.4%, 0.96, 0.89, 0.92
A063  [135] 2019 Angiographic CT MRI CE-Net Lungs, Vessels, Eyes Sensitivity, Accuracy, AUC 0.8309, 0.9545, 0.9779
A064  [136] 2021 Colonoscopy ColonSegNet Polyp AP, IoU, DSC, Jaccard, Precision, Recall 0.8000, 0.8100, 0.7239, 0.8435, 0.8496, 0.9493
It has been used for the segmentation of dermatology images to detect lesions in the skin [111, 112], for the segmentation of PET images to detection tumour in human lungs [113], and for the analysis of microscopic images [115]. Few studies improved mask R-CNN with some modifications. For example, a dual mask R-CNN is used for the segmentation of PET and CT images to detect tumour and prostrate [114]. Another enhanced version of mask R-CNN called URCNN [116] is used for the segmentation of microscopic images to detect anthrax and some other tissue diseases. Similarly, another variant is a faster R-CNN [117] which is used for the segmentation of microscopic images to detect cancer cells. Further R-CNN’s improved variants include cascade R-CNN [120] used for the segmentation of ultrasound images to detect thyroid nodules, another variant is used for the segmentation of breast ultrasound images to detect lesions [119], a hybrid model of R-CNN and pyramid network (FPN) for the detection and segmentation of different multi-class artifacts in the endoscopy images [121], and a cascade R-CNN is used for the detection and segmentation of artefacts in the endoscopic images [122].
@@
FPN has also been improve to improve varients including Fourier feature pyramid network (FFPN) used for for the segmentation of ultrasound images [123], FedNet for the segmentation of polyp in the endoscopic images [124], and FPN with other CNN networks for the detection and segmentation of polyp in the endoscopy images  [125]. Su et al.  [126], suggested an improved version of CNN named fCNN standing for fast scanning deep convolutional neural network for the segmentation and identification of regions in histopathology images of breast. A recent article [118] has tested three two-stages model namely cascade R-CNN, faster R-CNN, and double head R-CNN on ultrasound images of arteries for their segmentation to detect various abnormalities. Another study a multi-task segmentation approach for the segmentation of retina for the detection of diabetic retinopathy [127]. Similarly, the fully connected convolutional neural networks have been used for the segmentation of retinal images [128, 129], for the segmentation of PET images  [132, 131] and CT images to detect tumour [131]. Another approach is the deep convolutional neural network for the segmentation of biopsy images to detect dermis, epidermis, and non-tissue regions [130].
@@
Double U-Net [133] is an improved version of the standard U-Net in which two U-Nets are stacked on the top of each other. The model was evaluated with four publicly available dataset of different modalities e.g., dermoscopy, colonoscopy, and microscopy [133]. Attention U-Net and deep convolutional neural network have been used for the segmentation and classification of sonography ultrasound images for the detection of breast cancer [134]. CE-Net (context encoder network) [135] is another model which is applied for the segmentation of 2D medical imaging. Similarly, ColonSegnet [136] is used for the detection, localization, and segmentation of polyp in the colonoscopy images.
@@ 5.2 Articles Related to One-Stage Detection
One-stage object detection models used for the segmentation of biomedical images include YOLO, SSD, YOLOv1/v2/v3/v4, RefineNet, and RetinaNet. In this section, these models, as applied to biomedical image segmentation, are discussed in detail along with other relevant considerations. In this regard, YOLO [137] has been used to localize and detect various active organs in 3D PET imaging modalities. YOLOv3 has been applied for the segmentation of MRI images to detect herniated intervertebral discs [54], and for the segmentation of 2D and 3D CT images to localize kidneys and identify conditions such as cystic kidneys, hypertrophied kidneys, and tumoral kidneys [138]. Other studies have utilized YOLO [51] and YOLOv5 [139] for the segmentation of brain MRI images to detect tumors. Similarly, YOLOv4 has been used for the segmentation of ultrasound images to detect various organs [140], YOLOv7 for the segmentation of echocardiographic images to identify and segment the anatomical structures of the left heart [141], and YOLOv8 for processing X-ray angiography to segment the coronary artery [55]. Furthermore, the scaled-YOLOv4 has been applied for the detection of pulmonary embolism in CT angiograms [53], and YOLOv5 [57] has been used for multi-type lesion detection in angiography images. Santos et al.  [142] used YOLOV5 for lesion detection in fundus images. In another study, YOLOv8 model is used for the segmentation of angiography images to identify and detect coronary artery stenosis [56]. Pkhrel et al.  [143] used YOLOv8 along with the augmentation for the segmentation of coronary artery in the angiograms to detect diseases.
Some researchers have combined YOLO with other models to enhance performance. For instance, Chang et al.[52] employed YOLO alongside a fully convolutional neural network for the segmentation and classification of heart MRI images. YOLO, in combination with the GrabCut model, has been used for the segmentation of dermoscopic images to detect skin lesions[144]. Moreno et al. [145] applied a combination of two convolutional neural networks for the segmentation of the left ventricle in heart MRI images. Other hybrid models include the YOLOv2 and Fourier Photography Microscopy (FPM) in combination for the detection of WBC  [146]. The Faster R-CNN and YOLO are combined and used for the segmentation of chest X-Ray images to detect different abnormalities [147]. YOLO-RF [148] is a hybrid model of YOLO and random forest which has been used for diabetic detection.
@@
YOLOv3-arch [149] is another improvement of the YOLOv3 which has been used for the segmentation of CT images to identify cholelithiasis and classify gallstones. Zhuang et al.  [150] used YOLOv3 for the segmentation of left ventricle in the echocardiography. Similarly, Zhang et al.  [151] suggested the use of and enhanced model YOLOV3 one-stage segmentation model for segmentation of ultrasound images to detect the location of thyroid nodules. Similarly, YOLO-CSE [152] is an improvement of YOLOV5 which is used for the segmentation of muscle ultrasound to detect myositis. YOLOX [153] is used for the segmentation of ultrasound images to examine the fetus for congenital heart disease. Similarly, Ghose et al.  [154] suggested the use of an enhanced version of YOLOV5 to segment the colonoscopic images for the detection of polyp. YOLOV4-tiny [155] has been used for the segmentation of endoscopic images to detect polyp. Further studies applied YOLOV4 for the segmentation of endoscopic images to detect multi-class artefact [156], and YOLOV2 for the detection of breast cancer in histopathological breast images [157]. YOLO has been for the segmentation of nucleus and characterization of tumour microenvironment [39], and for detection of nucleus and U-Net for the segmentation of nucleus in fluorescence microscopy [158]. YOLO has also been used for the identification of region proposal networks to diagnose cancer in breast histopathology images [159].
Durak et al.  [160] implemented different types of deep learning models for the detection of polyp in the gastro endoscopic images. Other different deep learning models including mask R-CNN, U-Net, CNN, and YOLO have been used for the segmentation of nucleus in the histopathological images [161]. Similarly, CNN combined with YOLO has been used for the detection of glaucoma in the color fundus images [162]. Nair et al.  [163] applied YOLOV4 for the nucleus detection in the histopathology images of breast. Similarly, Pal et al.  [164] applied YOLOV3 for detection of lesions in the retinal fundus images.
@@
Other articles demonstrate the use of different models for more diseases. For example, YOLOv3 has been used for the localization of the optic disc in retinopathy fundus images [165], YOLO for lesion detection in skin using dermoscopic images [166], and fuzzy logic combined with object detection models for cancer detection in dermoscopic images [167]. Additionally, the YOLOv3 network has been applied for the segmentation of PET/CT images to detect tumors in lymphoma patients [168]. Shwetha et al.[169] performed a two-stage approach for the detection of pus cells and epithelial cells. Similarly, Kong and Shen[170] proposed an enhanced version of YOLO for the segmentation of microscopic images to detect microorganisms. Another study applied a combined model of YOLO and U-Net for the segmentation of microscopic images to characterize the neuron [171].
@@
Furthermore, SSD and other segmentation models have been used for the detection of benign and malignant breast cancer in ultrasound images [60]. Similarly, SSD has also been applied for the detection of various lung abnormalities in ultrasound images [59], and for the detection of polyps in gastrointestinal endoscopic images using segmentation [62]. In a hybrid approach, Xu et al. [58] proposed DAD-SSD (Dynamic Attention Deconvolutional Single Shot Detector) for the classification of different types of polyps in colonoscopic images. MP-FSSD (Multi-scale Pyramidal Fusion Single Shot Multibox Detector Network) [61] is recently proposed as a modified form of single shot detector which has been used for for the segmentation of endoscopic images frames or the endoscopy to detect polyp.